# Discrete Solution Operator Learning for Geometry-Dependent PDEs

### v2

@@ 1 Introduction
This observation highlights a distinction that is often overlooked: although classical discrete numerical solvers can be regarded as operators, learning a continuous operator is not equivalent to learning a discrete numerical solution procedure. Classical solvers do not approximate a single global mapping; instead, they execute a sequence of fixed procedures—local contribution evaluation, global assembly, and solution—which is invariant under geometry variation, even though the assembled global information changes (Figs. 1a,b) [20, 27, 6, 4]. Approximating a smooth functional mapping may therefore be misaligned with the procedural nature of geometry-dependent PDE solution processes.
@@ 2.1 Geometry-dependent Poisson problem
We first consider a geometry-dependent elliptic problem as a minimal setting to isolate the effect of domain variation. In this case, the governing equation is fixed, while the computational domain and boundary configuration vary across samples. This setup allows us to examine whether an operator can generalize across discrete changes in domain topology and boundary activation, rather than merely interpolating solution patterns on a fixed geometry. The underlying equation is a scalar Poisson-type problem; detailed formulations are provided in the Supplementary Information A.1.
@@
Unlike standard operator-learning benchmarks defined on regular domains, geometry, boundary specifications, and source distributions are explicitly provided as discrete input channels (Fig. 2a), and the solution pattern is predicted on the same computational grid.
Although both inputs and outputs are represented on a regular grid, the effective computational domain and boundary conditions may vary substantially, including the presence of internal holes, non-smooth boundaries, and highly localized boundary segments. Under this formulation, the dominant source of problem complexity no longer arises from source terms or material parameters, but instead stems from geometry-induced changes in the discrete computational structure.
The training dataset consists of topologically simple, simply connected geometries with smooth outer boundaries. Boundary conditions are applied continuously along the external boundary, and source terms are generated as linear combinations of low-frequency modes. As a result, the corresponding reference solutions are globally smooth, making this a relatively mild benchmark among geometry-dependent PDE problems.
Despite its relative simplicity, this setting already exposes a fundamental distinction between continuous and discrete operator learning paradigms. Because geometric variations directly alter the set of active computational regions and boundary segments on the grid, successful generalization requires adapting to discrete changes in local contributions rather than smooth interpolation of solution patterns.
@@
Using identical training strategies and comparable model capacities, we train DiSOL, DeepONet, and FNO, and compare their convergence behavior and generalization performance. As shown in Fig. 2b, DiSOL converges rapidly during training and achieves substantially lower errors on both the training and validation sets. In contrast, the training error of FNO saturates at a relatively high level and does not decrease further with continued optimization. Under this setting, DeepONet attains lower validation errors than FNO, indicating relatively stronger fitting capacity. For conservative ID comparisons, we therefore adopt DeepONet as the primary continuous-operator baseline, while still reporting FNO results for completeness.
Figure 2c presents two representative ID test cases (more results are presented in Supplementary Information E.1). Since these samples follow the training distribution and yield globally smooth Poisson solutions, both continuous-operator baselines are able to capture the coarse solution trend. Nevertheless, compared with DiSOL, DeepONet (and FNO) exhibit consistent residual discrepancies in the normalized pattern, which are most visible near irregular geometric features and boundary-adjacent regions, where the solution is jointly shaped by the discrete activation of boundary segments and the local domain geometry. These deviations manifest as mild distortions in the relative distribution of values and weakened boundary-conditioned contrast, rather than random noise. DiSOL more faithfully reproduces both the global morphology and the boundary-modulated local variations, leading to noticeably lower errors under the same ID setting. A capacity-matched U-Net control study is provided in Supplementary Information H to disentangle backbone effects from the discrete-operator formulation.
We further assess generalization under strongly OOD conditions (Figs. 2d and S6), where multiple unseen factors are introduced simultaneously. These include geometries with sharp corners and internal holes, which induce topological changes and newly formed internal boundaries, spatially discontinuous and highly localized boundary conditions, and source terms containing substantial high-frequency components. Despite these challenges, DiSOL remains stable and produces solution patterns that closely match the reference solutions in both overall structure and key local features.
In contrast, DeepONet exhibits pronounced structural failure under OOD geometric conditions. Its predictions not only suffer from localized errors, but also deviate from the reference solutions at the level of global solution patterns. This behavior indicates that, for geometry-dependent discrete problems, continuous operator paradigms struggle to robustly capture the discrete structural variations induced by significant geometric changes.
@@
Overall, even for the classical two-dimensional Poisson equation, where solution patterns are globally smooth and the underlying physical mechanisms are relatively simple, continuous neural operator methods already exhibit intrinsic limitations when the problem is formulated in a geometry-dependent and discrete manner. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL aligns more naturally with the principles of numerical solution procedures, thereby achieving robust and accurate predictions under both ID and OOD geometric conditions.
@@ 2.2 Geometry-dependent advection-diffusion problem
We next consider a steady advection-diffusion problem to examine geometry generalization under transport-dominated dynamics. Although the governing equation and physical parameters are fixed within each case, variations in domain geometry induce sharp changes in effective transport pathways and boundary-layer structures. This case therefore probes whether an operator can adapt to geometry-induced changes in solution structure when global transport effects play a dominant role. The precise formulation and parameter settings are provided in Supplementary Information A.2.
We adopt the same discrete input-output representation as in the Poisson case (Fig. 2a), consisting of three input channels: geometry mask, boundary-condition locations, and source field; and a single output channel representing the scalar solution pattern. In this problem, the Péclet number ($\mathrm{Pe}$) characterizes the relative strength of advection to diffusion [13], and separate operators are trained for two representative cases ($\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$). Under this design, increased difficulty arises from the interaction between transport-driven solution structures and geometry-induced discreteness, rather than from changes in the input encoding. It is worth emphasizing that, although the input representation is consistent across different PDEs, the advection-diffusion operator is inherently more complex: its dominant solution patterns are jointly shaped by transport direction, boundary conditions, and geometric constraints, rather than being determined solely by the source distribution.
Figures 3a and b present representative ID validation results for the diffusion-dominated ($\mathrm{Pe}\approx 0.45$) and advection-dominated ($\mathrm{Pe}\approx 4.5$) cases (more results are presented in Supplementary Information E.2). At lower Péclet numbers, the solution patterns remain relatively smooth but already exhibit asymmetric structures aligned with the transport direction. As observed, DiSOL accurately predicts the global solution patterns with their randomly selected boundaries. In contrast, FNO struggles to maintain structural consistency on irregular geometries, producing predictions with noticeable morphological distortions across multiple regions.
As the Péclet number increases, advection becomes the dominant mechanism governing solution behavior. The resulting solution structures exhibit stronger directional transport features and heightened sensitivity to geometric boundaries and local structural variations. Under this advection-dominated setting (Fig. 3b), DiSOL continues to capture the primary transport pathways and their coupling with complex geometries, yielding solution patterns that remain consistent with the reference numerical solutions in both global morphology and key local features. By contrast, FNO displays pronounced structural distortions, with predicted patterns failing to reflect the true transport behavior, particularly near geometric boundaries.
To further examine optimization behavior under varying transport conditions, Figs. 3c and d show the training and validation loss histories for the two Péclet cases. In this transport-dominated setting, FNO serves as the primary continuous-operator baseline, while DeepONet results are also reported for completeness. For both $\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$, DiSOL exhibits stable convergence and achieves substantially lower training and validation errors. In contrast, continuous neural operator baselines show early error saturation in both cases, and their relative convergence behavior does not fundamentally improve with increasing transport dominance. This observation suggests that transport strength itself is not the primary source of difficulty in operator learning.
@@
We further evaluate generalization under strongly OOD geometric conditions at high Péclet numbers (Figs. 3e, f and S10), where geometries with sharp corners and topological changes are combined with highly localized and spatially discontinuous boundary conditions. Although these settings substantially amplify geometry-induced discrete structural variations, DiSOL maintains stable predictive performance, with error distributions that are statistically superior to those of continuous operator baselines. The corresponding boxplot analysis of relative $L_{1}$ errors indicates that performance degradation is primarily driven by geometric distribution shifts, rather than by the increase in the Péclet number itself.
Taken together, these results demonstrate that, for the advection-diffusion equation governed jointly by geometry and boundary conditions, the limitations of continuous neural operator methods do not stem from increased transport strength alone. Instead, they arise because transport-dominated dynamics are highly sensitive to geometry-induced discrete changes, which can markedly alter the solution. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL more naturally captures the coupling between geometry and transport effects, thereby achieving enhanced stability and generalization performance in both ID and OOD settings.
@@ 3 Discussion
A central insight emerging from our results is that the primary challenge in geometry-dependent PDE learning does not stem from model capacity or optimization, but from the nature of geometric variation itself. Importantly, this behavior should not be interpreted as a fundamental limitation of continuous neural operator learning. Specifically, in many practical settings, changes in geometry induce discrete and non-smooth structural variations—such as alterations in domain topology, boundary classification, or active computational regions—that fundamentally modify how local contributions are assembled. When such geometry-induced effects dominate problem complexity, the mapping from problem specification to solution field is no longer well described as interpolation within a smooth function-space operator. Continuous operator methods remain highly effective when variations in problem specification, including geometry, can be treated as smooth perturbations of an underlying function-space mapping.
However, geometry-dependent problems often violate this smooth-variation premise. While the governing physical laws and numerical solution procedures remain invariant, geometric changes discretely alter how these procedures are instantiated on a computational domain. In this sense, what changes is not the underlying algorithmic logic, but the discrete structure through which local contributions are activated, combined, and propagated. Within this perspective, the advantage of DiSOL does not arise from a particular network architecture, but from its alignment with the discrete nature of geometry-dependent numerical problems. By operating directly on discretized representations, the learned operator mirrors the structure of classical numerical solvers, explicitly accommodating geometry-induced structural variation without forcing such effects into a smooth function-space interpolation framework.
@@ 5 Acknowledgements
Support from the National Natural Science Foundation of China (Grant nos. T2488101 and 12032014) is acknowledged (X.-Q. Feng and J. Bai). Support from the National Natural Science Foundation of China (Grant no. 12502234) is acknowledged (J. Bai).

### v3

@@ 1 Introduction
This observation highlights a distinction that is often overlooked: although classical discrete numerical solvers can be regarded as operators, learning a continuous operator is not equivalent to learning a discrete numerical solution procedure. Classical solvers do not approximate a single global mapping; instead, they execute a sequence of fixed procedures (local contribution evaluation, global assembly, and solution) which is invariant under geometry variation, even though the assembled global information changes (Figs. 1a,b) [21, 28, 7, 5]. Approximating a smooth functional mapping may therefore be misaligned with the procedural nature of geometry-dependent PDE solution processes.
@@ 2.1 Geometry-dependent Poisson problem
We first consider a geometry-dependent elliptic problem as a minimal setting to isolate the effect of domain variation. The governing equation is fixed, while the computational domain and boundary configuration vary across samples, including discrete changes in topology and boundary activation. We study a scalar Poisson-type problem; detailed formulations are provided in Supplementary Information A.1.
@@
Unlike standard operator-learning benchmarks defined on regular domains, geometry, boundary specifications, and source distributions are provided as discrete input channels (Fig. 2a), and the solution field is predicted on the same embedded grid. Although the representation is grid-based, geometric variation changes the effective computational region and boundary segments, so generalization requires adapting to discrete structural changes rather than smooth interpolation on a fixed domain.
The training dataset contains simply connected geometries with smooth outer boundaries, continuous boundary conditions along the exterior boundary, and low-frequency source terms, yielding globally smooth reference solutions. Even in this mild regime, geometry-induced changes in active regions and boundary segments expose a fundamental distinction between continuous and discrete operator learning.
@@
Using identical training strategies and comparable model capacities, we train DiSOL, DeepONet, and FNO. As shown in Fig. 2b, DiSOL converges rapidly and achieves substantially lower training/validation errors, whereas FNO saturates early; DeepONet provides a stronger continuous-operator baseline in this setting and is used as the primary continuous baseline (FNO is reported for completeness).
Figure 2c shows representative ID test cases (additional examples in Supplementary Information E.1). While DeepONet/FNO capture the coarse trend, they exhibit systematic residual distortions near irregular geometric features and boundary-adjacent regions, leading to higher errors than DiSOL. DiSOL more faithfully reproduces both global structure and boundary-modulated local variations under the same ID setting. A capacity-matched U-Net control study is provided in Supplementary Information H.
We additionally benchmark three representative geometry-aware variants of continuous neural operators, including Diffeomorphic Mapping Operator Learning (DIMON) [36], Geo-FNO [17], and GNO [19]. They implement different geometric inductive bias via explicit geometric representations or graph-based aggregation. These variants yield at most modest improvements over vanilla DeepONet/FNO and remain clearly below DiSOL under both ID and geometry OOD shifts; detailed comparisons are provided in Supplementary Information E.1.
We further assess generalization under strongly OOD conditions (Figs. 2d and S6) that combine sharp corners and internal holes (topology changes), highly localized/discontinuous boundary conditions, and higher-frequency source components. DiSOL remains stable and closely matches the reference solutions, whereas DeepONet exhibits pronounced failures, including deviations in global solution patterns rather than only localized errors.
@@
Overall, even for the classical two-dimensional Poisson equation with globally smooth solutions, continuous neural operator methods show intrinsic limitations when geometry induces discrete structural variation in the instantiated computational domain. By learning and assembling local contributions directly on discretized domains, DiSOL aligns with numerical solution procedures and achieves robust predictions under both ID and OOD geometric conditions.
@@ 2.2 Geometry-dependent advection-diffusion problem
We next consider a steady advection-diffusion problem to examine geometry generalization under transport-dominated dynamics. While the governing equation and physical parameters are fixed within each Péclet case, variations in domain geometry alter effective transport pathways and boundary-layer structures, providing a stringent test of generalization under geometry-induced structural change. The formulation and parameter settings are provided in Supplementary Information A.2.
We use the same embedded-grid representation as in the Poisson case (mask, boundary-condition map, and source as inputs; scalar field as output) and train separate operators for two regimes, $\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$. Representative ID validation examples are shown in Figs. 3a–b (additional results in Supplementary Information E.2). Across both regimes, DiSOL captures the main transport structures and their coupling with complex geometries, whereas continuous-operator baselines exhibit noticeable structural distortions, particularly near geometric boundaries and irregular features.
Figures 3c–d report training/validation histories for the two Péclet cases. DiSOL converges stably and reaches substantially lower errors, while continuous baselines saturate early in both regimes, suggesting that increased transport dominance alone is not the primary source of difficulty in operator learning under geometry variation (DeepONet is reported for completeness).
@@
We then evaluate generalization under strongly OOD geometries at high Péclet numbers (Figs. 3e–f and S10), combining sharp corners/topology changes with highly localized and spatially discontinuous boundary conditions. DiSOL maintains stable performance with superior error distributions, and the relative $L_{1}$ boxplots indicate that degradation is driven primarily by geometry shifts rather than the increase in $\mathrm{Pe}$.
As an additional test, we benchmark geometry-aware variants of continuous neural operators (GNO and Geo-FNO, and DIMON where applicable). These approaches provide at most modest gains over vanilla DeepONet/FNO and remain clearly below DiSOL under both ID and geometry OOD shifts; full results are reported in Supplementary Information E.2.
Taken together, these results show that the limitations of continuous neural operator baselines are not explained by transport strength alone, but by geometry-induced discrete structural changes that strongly modulate transport-dominated solutions. By learning and assembling local contributions directly on discretized domains, DiSOL captures the coupling between geometry and transport effects and achieves robust generalization in both ID and OOD settings.
@@ 3 Discussion
A central insight emerging from our results is that the primary challenge in geometry-dependent PDE learning does not stem from model capacity or optimization, but from the nature of geometric variation itself. Importantly, the degradation of continuous-operator baselines should not be interpreted as a fundamental limitation of continuous neural operator learning. Specifically, in many practical settings, changes in geometry induce discrete and non-smooth structural variations—such as alterations in domain topology, boundary classification, or active computational regions—that fundamentally modify how local contributions are assembled. When such geometry-induced effects dominate problem complexity, the mapping from problem specification to solution field is no longer well described as interpolation within a smooth function-space operator. Continuous operator methods remain highly effective when variations in problem specification, including geometry, can be treated as smooth perturbations of an underlying function-space mapping. This interpretation is further supported by additional benchmarks against representative geometry-aware variants of continuous neural operators, including GNO, Geo-FNO, and DIMON. These methods bring geometric inductive bias through explicit geometric representations or mapping/alignment mechanisms, and can yield modest improvements over vanilla DeepONet/FNO. However, our results show that they do not eliminate the dominant failure modes under strong geometry shifts (see Supplementary Information E), indicating that geometric awareness alone is insufficient when geometry induces discrete structural changes in the instantiated solver structure.
While the governing physical laws and numerical solution procedures remain invariant, geometric changes discretely alter how these procedures are instantiated on a computational domain. In this sense, what changes is not the underlying algorithmic logic, but the discrete structure through which local contributions are activated, combined, and propagated. Within this perspective, the advantage of DiSOL does not arise from a particular network architecture, but from its alignment with the discrete nature of geometry-dependent numerical problems. By operating directly on discretized representations, the learned operator mirrors the structure of classical numerical solvers, explicitly accommodating geometry-induced structural variation without forcing such effects into a smooth function-space interpolation framework.
@@ 4.4 Training objective and evaluation metrics
5 Data availability
All datasets used in this study are publicly available on Zenodo (DOI: https://doi.org/10.5281/zenodo.18639634) [4], including the full datasets for all four benchmark cases.
6 Code availability
The full source code for DiSOL is available at https://github.com/JinshuaiBai/Discrete-Solution-Operator-Learning-DiSOL. An executable Code Ocean compute capsule is also provided for reproducibility and ease of use at https://codeocean.com/capsule/4396603/tree/v1.
@@ 7 Acknowledgements
Support from the National Natural Science Foundation of China (Grant no. T2488101) is acknowledged (X.-Q. Feng and J. Bai). Support from the National Natural Science Foundation of China (Grant no. 12502234) is acknowledged (J. Bai).